OpenStudy (anonymous):

. Let x and y be two positive numbers such the sum is 175 and (x+3) (y+4) is a maximum.

OpenStudy (anonymous):

since \(x+y=175\) you have \(y=175-x\) maximize \((x+3)(175-x+4)\) after doing some algebra first

OpenStudy (anonymous):

Oh so you don't start with the (x+3) and (y+4)= 175?

OpenStudy (anonymous):

So you first find x then input X to maximize?

OpenStudy (anonymous):

it says "the sum is 175" which translates as \(x+y=175\)

OpenStudy (anonymous):

yeah find the value of \(x\) that gives the max it is a quadratic, so the max is at the vertex

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

Yeah so I found X= 175/2. In order to find Y. I must input x to the equation. (175/2 + 3) (y + 4) = 175?

OpenStudy (anonymous):

some mistake there i think

OpenStudy (anonymous):

\[(x+3)(175-x+4)=(x+3)(179-x)=-x^2+176 x+537\]

OpenStudy (anonymous):

then vertex is at \(\frac{176}{2}=88\)

OpenStudy (anonymous):

Does this relate to Optimization because I taught I was supposed to find a x and y coordinate?

OpenStudy (anonymous):

you have \(x=88\) and since \(x+y=175\) you can certainly find \(y\) and also you can find \((x+3)(y+4)\)

OpenStudy (anonymous):

Oh alright now I get it thanks.

OpenStudy (anonymous):

yw